OpenStudy (anonymous):

whats -13/2 and -7/2?

OpenStudy (anonymous):

-13/2=-6.5 and -7/2=-3.5

hero (hero):

If you want to use mix numbers then: \[-\frac{13}{2} = -\frac{12}{2} - \frac{1}{2} = -\left(\frac{12}{2} + \frac{1}{2} \right) = -\left(6 + \frac{1}{2} \right) = -6 \frac{1}{2}\]

hero (hero):

You can do the same for the other fraction